OpenStudy (anonymous):

ADD.8 over 8+z + 4 over 8+z= wouldn't it be 12+z

OpenStudy (anonymous):

I came up with 12/z+8

OpenStudy (anonymous):

Unless I"m reading it wrong

OpenStudy (anonymous):

Thats what I got and its correct

OpenStudy (anonymous):

I am just trying to correct the ones I got wrong on my exam

OpenStudy (anonymous):

\[\frac{8}{8+z}+\frac{4}{8+z}\] \[=\frac{12}{8+z}\]